# Nested, asymmetric H–He circumstellar shells in the Type Icn/Ibn SN 2024abvb

### v1

@@ 1 Introduction
Understanding the evolution and the ultimate fate of stars remains one of the key questions in astrophysics. The INteracting Transients at ESO-vLt (INTEL) program aims to investigate the role of circumstellar medium (CSM) interaction in various species of stellar transients. Targets of the INTEL program are stellar transients with observational evidence of slowly expanding CSM, such as tidal-disruption events (TDEs; e.g., Gezari, 2021), interacting gap transients (Pastorello & Fraser, 2019), interacting SNe (Smith, 2014; Fraser, 2020) and super-luminous SNe (SLSNe; Gal-Yam, 2019; Inserra, 2019). This program uses the Very Large Telescope (VLT) Ultraviolet and Visual Echelle Spectrograph (UVES, Dekker et al., 2000) and the X-Shooter spectrograph (Vernet et al., 2011) for high-resolution spectroscopy, while the FOcal Reducer/low dispersion Spectrograph 2 (FORS2, Appenzeller et al., 1998) for polarimetric data. This science programme operates in parallel with low-resolution spectroscopy efforts and is intended to complement and enhance them, with a distinct and differing scientific focus.
@@
To complicate the overall picture, SNe with intermediate properties between two different types are occasionally observed, including the transitional SNe IIn/Ibn (whose spectra show narrow H and He I with comparable strengths; Pastorello et al., 2008b, 2015a; Smith et al., 2012b; Reguitti et al., 2022; Farias et al., 2024; Gangopadhyay et al., 2025) and SNe Ibn/Icn (with spectra showing at the same time He I and C/N/O lines; e.g., Pursiainen et al., 2023). The recent discovery of the nearby SN 2024abvb provides a rare opportunity to study in detail a SN with transitioning properties between those of a Type Icn and a Type Ibn SN.
@@ 2 Observations
SN 2024abvb (Hu et al., 2026, Aster et al. submitted) was discovered by the Asteroid Terrestrial-impact Last Alert System (ATLAS, Smith et al., 2020) on 2024 November 22 (Tonry et al., 2024) and subsequently classified by the Nordic-optical-telescope Un-biased Transient Survey (NUTS, Mattila et al., 2016) on 2024 November 28 as a Type Icn SN (Stritzinger et al., 2024) at z = 0.039 (see Aster et al. 2026 submitted), equivalent to a luminosity distance of 164.8 Mpc using the following cosmology measurements $H_{0}=73.04$ km s-1 Mpc-1, $\Omega_{M}=0.326$ and a flat Universe (Riess et al., 2022). The Galactic reddening toward the SN position is E(B-V) = 0.164 mag (Schlafly & Finkbeiner, 2011). None of our high-resolution spectra show Na id lines related to host galaxy reddening, hence we assume the Galactic reddening as the total. This is also consistent with the remote location of SN 2024abvb (See Section 2.2). The light curve of SN 2024abvb from publicly available feeder surveys data [^1] https://www.wis-tns.org/object/2024abvb is shown in Fig. 1 with respect to the maximum light (MJD 60645.13, Aster et al. 2026 submitted and references therein) to guide the spectroscopic temporal analysis.
@@ 2.2 The host and environment of SN 2024abvb
The host galaxy of SN 2024abvb, WISEA J011055.76-054416.6, has a magnitude of $g=17.3$, and its nucleus lies 28″ from the location of the SN. It appears to be a fairly normal spiral galaxy, close to face-on to our line of sight. The galaxy has the same redshift of the SN ($z=0.039$) as shown by the low-resolution host galaxy spectrum (see Aster et al. 2026 submitted), and hence has an absolute magnitude of g=18.9g=-18.9, while the projected separation from the SN is 21.5 kpc in the plane of sky alone.
@@
We compare the projected host-galaxy offsets of SN 2024abvb with the distributions of SNe Ib, Ic, IIn, and Ibn from the PTF sample (Schulze et al., 2021). Fig. 3 shows the cumulative distributions, demonstrating that our event lies in the extreme tail of the offset distribution, particularly for SNe Ibn. We note that there are insufficient data to perform a similar analysis for SNe Icn. Schulze et al. (2021) report that the majority of SNe (92%) fall within the 80% light radius ($r_{80}$), which, together with galaxy mass, traces the stellar mass content of galaxies independently of type and star formation activity (Miller et al., 2019; Mowla et al., 2019). A first-order approximation for spiral galaxies gives $r_{e}/r_{80}\approx 1$ (Graham, 2001), implying that SN 2024abvb is indeed located in an extreme region. Aster et al. 2026 (submitted) report a host-galaxy mass of $\log(M/M_{\odot})=10.33^{+0.11}_{-0.10}$ for SN 2024abvb, that combined with the SN position, it implies that the event lies outside the $r_{80}$ radius.
@@ 3.1.1 Pre peak UVES and FORS2 spectra
Our FORS2 and UVES spectra were taken at 2-2 days from maximum light. While the former covers a wider wavelength range, it does not allow us to resolve the narrow features seen in the UVES spectrum of SN 2024abvb (Figs. 16 and 17). They both show a hot blue continuum. We cannot accurately measure the blackbody temperature as in the subsequent XShooter epochs because the peak of the first spectrum is to the blue of 3000 Å, while in the second epoch, we are unable to fit both the blue and red parts of the spectrum, possibly due to the effect of absorption by metals. Most of the features bluer than 4500 Å observed in the FORS2 spectrum and the ALFOSC classification spectrum can be attributed to O ii and Ne ii, while no lines corresponding to higher ionisation states of carbon are detected, unlike what is observed in SNe Icn (e.g. Pellegrino et al., 2022; Nagao et al., 2023).
@@
The two strongest emission features in the FORS2 spectrum, also visible in the ALFOSC classification spectrum (Fig. 4), are at 5890 Å and 7236 Å, and we associate these with C ii. While the lines exhibit a broad base (more evident in the UVES data; see, e.g. in Fig. 17), we do not attribute this to ejecta. The wings of the broad feature extend out to 2000 km s-1, and these likely arise from electron scattering in the CSM (e.g. Chugai, 2001; Fransson et al., 2014). This implies an electron temperature of $T_{e}\sim 10^{4}$ K and an electron-scattering optical depth of order unity. This requires a compact, highly ionised CSM with electron densities of $n_{e}\sim 10^{9}$$10^{10}$ cm-3 at radii $R\sim 10^{14}$$10^{15}$ cm, roughly matching what inferred from the bolometric luminosity (see Alster et al 2026, submitted; Hu et al., 2026). Such conditions are inconsistent with a steady stellar wind and instead point to eruptive or binary-driven mass-loss episodes prior to explosion.
@@
We also observe a line profile around 6578 Å where another C ii line, expected to have a strength comparable to that of $\lambda$7236 due to its role in the recombination cascade, should be present. In addition, the C ii $\lambda$4267 line, which shares the same lower energy level as $\lambda$5890 but has an Einstein coefficient for spontaneous emission ($A_{ki}$) approximately an order of magnitude higher, is also detected in the FORS2 and ALFOSC spectra, albeit with a P Cygni profile (assuming that the prominent blue-shifted absorption is indeed associated with C ii).
@@
In the UVES spectrum, turning to the other lines, we see multiplets 15 and 16 of C ii at wavelengths $5640-5662\penalty 10000\ \AA$ and $5133-5151\penalty 10000\ \AA$ , respectively. Although these are at relatively low S/N, we measure a velocity from the P-Cygni absorption component minima of these lines of $\sim$100 km s-1. We also see strong lines of O ii, including multiplets 1, 9, 24, 25, 28, and 32. Multiplet 1 is particularly strong, with numerous lines between 4639 and 4696 Å. Al ii appears to be present with multiplets 1 and 2.
@@
Besides the weak P-Cygni absorptions, two lines in emission stand out at 5889 Å and 6003 Å. The bluer line is also shown in the classification spectrum and, as mentioned above, it is consistent with C ii. However, due to the different profile displayed by this line with respect to other C ii lines, we cannot rule out that such a profile belongs to Na iD. The fact that only the emission component is observed would suggest that it arises in a Na-rich CSM shell, similarly to what is observed in the type IIn SN 2011A (de Jaeger et al., 2015), although in that case the Na iD displayed both an emission and blue-shifted absorption. With respect to the 6003Å line, we also checked for diffuse interstellar bands (DIBs) using the catalogue of Jenniskens & Desert (1994) with updates as listed online [^4] https://leonid.arc.nasa.gov/DIBcatalog.html, but find no convincing matches. Furthermore, no absorption resonance lines such as Na iD and Ca H&K have been observed in any of the spectra, suggesting that none of the observed lines could be associated with the ISM.
@@
Looking in detail at our high resolution UVES spectrum, multiple He i lines are seen, most clearly the strong $\lambda$5875 line which has a pronounced P-Cygni profile (see Fig. 17), with a velocity minimum at 120 km s-1, and a blue edge extending out to $\sim$190 km s-1, broadly consistent with the C ii. Such a line is comparable in strength to O i lines, especially those of multiplet 1. Other He lines are visible, including $\lambda$5015 and $\lambda$4921 (Fig. 16), but these lines are weaker and only the absorption component is observed. This is consistent with what is observed for He i $\lambda$5875, where the absorption is stronger than the emission component. The presence of narrow He lines, identifiable only thanks to our high resolution, would imply a Type Ibn classification for SN 2024abvb, although we note that the weakness of these features could also lead to labelling it as a transitional Ibn/Icn SN (analogous to the transitional Ibn/IIn SNe; Pastorello et al., 2015b). This suggests that an empirical classification based on low-resolution spectra may be unreliable, especially when interaction with a CSM is present.
@@ 3.1.2 XShooter post peak spectra
C ii lines are also prominently detected, with all expected strong transitions clearly visible. These lines, like He i, display P Cygni profiles with a velocity of roughly 120 km s-1 as measured from the absorption component. In contrast, there is no evidence for the presence of He ii, C iv, N ii or N i lines in the spectra. The presence of C iii and O iii remains uncertain; if present, these features are extremely weak and appear only in absorption. Nevertheless, these tentative identifications, along with the non-detection of higher ionisation lines, place constraints on the ionising radiation field, limiting it to a maximum of approximately 35–40 eV, with the bulk of ionisation radiation likely peaking in the 10–25 eV range (Moore, 1945). O i lines are also detected and are mostly attributable to recombination processes. [O i] $\lambda 6300$ is detected in emission at $\phi=17.2$d with a velocity width of the order of $\sim 100\penalty 10000\$km s-1, suggesting a CSM origin.
@@
Among other lines, Si ii is identified, with multiplets 1, 2, and 3 being present in the spectra. Multiplet 4 is absent, and the strongest line of multiplet 5, if present, appears extremely weak. Mg ii is observed, particularly in multiplets 4, 6, and possibly 8, which seem to appear solely in emission. Conversely, multiplets 2 and 10 are undetected, which is consistent with their known atomic parameters. Lines at 5696.6 and 5722.7 Å could be tentatively identified as Al iii (multiplet 2). Al ii (multiplets 1 and 2) is still detected as in the UVES spectrum. This identification excludes the possibility that the feature at 5696.0 Å is due to C iii (singlet transition). However, the line at 4325.7 Å remains unidentified, with multiple possible line assignments requiring further investigation. Notably, as in the FORS2 and classification spectra, none of the C iii triplet transitions are detected. Fe ii lines are observed prominently in the XShooter spectra, particularly from multiplet 42, as well as multiplets 49, 48, and 37. These lines show P Cygni profiles, at times complex due to proximity with other lines, displaying an absorption minimum velocity of 100 km s-1. By the time of the second XShooter spectrum ($\phi=17.2$d; Fig. 18), Fe ii features have significantly strengthened, suggesting the possible formation of an “iron blanket” in the near-ultraviolet region ($\leq 3500$Å) similar to the usual pseudo-continuum formed by Fe lines in SNe IIn (e.g. Smith, 2017, and references therein). In addition to the emission lines at 5889 Å and 6003 Å observed in the UVES spectrum, another emission line at 8235 Å is identified in the XShooter spectra. For this feature, potential associations include C ii, Mg ii and O i. However, since its profile differs from those of other C ii, Mg ii and O i lines, the precise nature remains uncertain. This feature may represent a blend of contributions from all three species.
@@
The ionisation potentials for Mg ii and Si ii, assuming the observed lines arise from recombination, are approximately 15 and 16 eV, respectively. Their detection thus broadens the inferred ionisation range while slightly lowering the average ionising energy with respect to the analysis of the non-metal elements. This reinforces the conclusion that the source is dominated by relatively soft UV radiation.

### v2

@@ 1 Introduction
Understanding the evolution and the ultimate fate of stars remains one of the key questions in astrophysics. The INteracting Transients at ESO-vLt (INTEL) program aims to investigate the role of circumstellar medium (CSM) interaction in various species of stellar transients. Targets of the INTEL program are stellar transients with observational evidence of slowly expanding CSM, such as tidal-disruption events (TDEs; e.g., Gezari, 2021), interacting gap transients (Pastorello & Fraser, 2019), interacting SNe (Smith, 2014; Fraser, 2020), SNe interacting with a confined CSM (also known as ‘flash features’, e.g. Bruch et al., 2023; Jacobson-Galán et al., 2024) and super-luminous SNe (SLSNe; Gal-Yam, 2019; Inserra, 2019). This program uses the Very Large Telescope (VLT) Ultraviolet and Visual Echelle Spectrograph (UVES, Dekker et al., 2000) and the X-Shooter spectrograph (Vernet et al., 2011) for high-resolution spectroscopy, while the FOcal Reducer/low dispersion Spectrograph 2 (FORS2, Appenzeller et al., 1998) for polarimetric data. This science programme operates in parallel with low-resolution spectroscopy efforts and is intended to complement and enhance them, with a distinct and differing scientific focus.
@@
To complicate the overall picture, SNe with intermediate properties between two different types are occasionally observed, including the transitional SNe IIn/Ibn (whose spectra show narrow H and He I with comparable strengths; Pastorello et al., 2008b, 2015; Smith et al., 2012b; Reguitti et al., 2022; Farias et al., 2024; Gangopadhyay et al., 2025) and SNe Ibn/Icn (with spectra showing at the same time He I and C/N/O lines; e.g., Pursiainen et al., 2023). The recent discovery of the nearby SN 2024abvb provides a rare opportunity to study in detail a SN with transitioning properties between those of a Type Icn and a Type Ibn SN.
@@ 2 Observations
SN 2024abvb (Aster et al., 2026; Hu et al., 2026; Shi et al., 2026) was discovered by the Asteroid Terrestrial-impact Last Alert System (ATLAS, Smith et al., 2020) on 2024 November 22 (Tonry et al., 2024) and subsequently classified by the Nordic-optical-telescope Un-biased Transient Survey (NUTS, Mattila et al., 2016) on 2024 November 28 as a Type Icn SN (Stritzinger et al., 2024) at z = 0.039 (see Aster et al., 2026), equivalent to a luminosity distance of 164.8 Mpc using the following cosmology measurements $H_{0}=73.04$ km s-1 Mpc-1, $\Omega_{M}=0.326$ and a flat Universe (Riess et al., 2022). The Galactic reddening toward the SN position is E(B-V) = 0.164 mag (Schlafly & Finkbeiner, 2011). None of our high-resolution spectra show Na id lines related to host galaxy reddening, hence we assume the Galactic reddening as the total. This is also consistent with the remote location of SN 2024abvb (See Section 2.2). The light curve of SN 2024abvb from publicly available ATLAS (Tonry et al., 2018), Pan-STARRS (Chambers et al., 2019), BlackGEM (Groot et al., 2024), GOTO (Steeghs et al., 2022) and ZTF (Bellm et al., 2019) surveys data on the Transient Name Server (TNS [^1] https://www.wis-tns.org/object/2024abvb)) is shown in Fig. 1 with respect to the maximum light (MJD 60645.13, Aster et al., 2026, and references therein) to guide the spectroscopic temporal analysis.
@@ 2.2 The host and environment of SN 2024abvb
The host galaxy of SN 2024abvb, WISEA J011055.76-054416.6, has a magnitude of $g=17.3$, and its nucleus lies 28″ from the location of the SN. It appears to be a fairly normal spiral galaxy, close to face-on to our line of sight. The galaxy has the same redshift of the SN ($z=0.039$) as shown by the low-resolution host galaxy spectrum (see Aster et al., 2026), and hence has an absolute magnitude of g=18.9g=-18.9, while the projected separation from the SN is 21.5 kpc in the plane of sky alone.
@@
We compare the projected host-galaxy offsets of SN 2024abvb with the distributions of SNe Ib, Ic, IIn, and Ibn from the PTF sample (Schulze et al., 2021). Fig. 3 shows the cumulative distributions, demonstrating that our event lies in the extreme tail of the offset distribution, particularly for SNe Ibn. We note that there are insufficient data to perform a similar analysis for SNe Icn. Schulze et al. (2021) report that the majority of SNe (92%) fall within the 80% light radius ($r_{80}$), which, together with galaxy mass, traces the stellar mass content of galaxies independently of type and star formation activity (Miller et al., 2019; Mowla et al., 2019). A first-order approximation for spiral galaxies gives $r_{e}/r_{80}\approx 1$ (Graham, 2001), implying that SN 2024abvb is indeed located in an extreme region. Aster et al. (2026) report a host-galaxy mass of $\log(M/M_{\odot})=10.33^{+0.11}_{-0.10}$ for SN 2024abvb, that combined with the SN position, it implies that the event lies outside the $r_{80}$ radius.
@@ 3.1.1 Pre peak UVES and FORS2 spectra
Our FORS2 and UVES spectra were taken at 2-2 days from maximum light. While the former covers a wider wavelength range, it does not allow us to resolve the narrow features seen in the UVES spectrum of SN 2024abvb (Figs. 19 and 20). They both show a hot blue continuum. We cannot accurately measure the blackbody temperature as in the subsequent XShooter epochs because the peak of the first spectrum is to the blue of 3000 Å, while in the second epoch, we are unable to fit both the blue and red parts of the spectrum, possibly due to the effect of absorption by metals. Most of the features bluer than 4500 Å observed in the FORS2 spectrum and the ALFOSC classification spectrum can be attributed to O $i$ and Ne $i$, while no lines corresponding to higher ionisation states of carbon are detected, unlike what is observed in SNe Icn (e.g. Pellegrino et al., 2022; Nagao et al., 2023).
@@
The two strongest emission features in the FORS2 spectrum, also visible in the ALFOSC classification spectrum (Fig. 4), are at 5890 Å and 7236 Å, and we associate these with C ii. While the lines exhibit a broad base (more evident in the UVES data; see, e.g. in Fig. 20), we do not attribute this to ejecta. The wings of the broad feature extend out to 2000 km s-1, and these likely arise from electron scattering in the CSM (e.g. Chugai, 2001; Fransson et al., 2014). This implies an electron temperature of $T_{e}\sim 10^{4}$ K and an electron-scattering optical depth of order unity. This requires a compact, highly ionised CSM with electron densities of $n_{e}\sim 10^{9}$$10^{10}$ cm-3 at radii $R\sim 10^{14}$$10^{15}$ cm, roughly matching what inferred from the bolometric luminosity (Aster et al., 2026; Hu et al., 2026). Such conditions are inconsistent with a steady stellar wind and instead point to eruptive or binary-driven mass-loss episodes prior to explosion.
@@
We also observe a line profile around 6578 Å where another C $i$ line, expected to have a strength comparable to that of $\lambda$7236 due to its role in the recombination cascade, should be present. In addition, the C $i$ $\lambda$4267 line, which shares the same lower energy level as $\lambda$5890 but has an Einstein coefficient for spontaneous emission ($A_{ki}$) approximately an order of magnitude higher, is also detected in the FORS2 and ALFOSC spectra, albeit with a P Cygni profile (assuming that the prominent blue-shifted absorption is indeed associated with C ii).
@@
In the UVES spectrum, turning to the other lines, we see multiplets 15 and 16 of C $i$ at wavelengths $5640-5662\penalty 10000\ \AA$ and $5133-5151\penalty 10000\ \AA$ , respectively. Although these are at relatively low S/N, we measure a velocity from the P-Cygni absorption component minima of these lines of $\sim$100 km s-1. We also see strong lines of O $i$, including multiplets 1, 9, 24, 25, 28, and 32. Multiplet 1 is particularly strong, with numerous lines between 4639 and 4696 Å. Al $i$ appears to be present with multiplets 1 and 2.
@@
Besides the weak P-Cygni absorptions, two lines in emission stand out at 5889 Å and 6003 Å. The bluer line is also shown in the classification spectrum and, as mentioned above, it is consistent with C ii. However, due to the different profile displayed by this line with respect to other C $i$ lines, we cannot rule out that such a profile belongs to Na iD. The fact that only the emission component is observed would suggest that it arises in a Na-rich CSM shell, similarly to what is observed in the type IIn SN 2011A (de Jaeger et al., 2015), although in that case the Na iD displayed both an emission and blue-shifted absorption. With respect to the 6003Å line, we also checked for diffuse interstellar bands (DIBs) using the catalogue of Jenniskens & Desert (1994) with updates as listed online [^4] https://leonid.arc.nasa.gov/DIBcatalog.html, but find no convincing matches. Furthermore, no absorption resonance lines such as Na iD and Ca H&K have been observed in any of the spectra, suggesting that none of the observed lines could be associated with the ISM.
@@
Looking in detail at our high resolution UVES spectrum, multiple He i lines are seen, most clearly the strong $\lambda$5875 line which has a pronounced P-Cygni profile (see Fig. 20), with a velocity minimum at 120 km s-1, and a blue edge extending out to $\sim$190 km s-1, broadly consistent with the C ii. Such a line is comparable in strength to O i lines, especially those of multiplet 1. Other He lines are visible, including $\lambda$5015 and $\lambda$4921 (Fig. 19), but these lines are weaker and only the absorption component is observed. This is consistent with what is observed for He i $\lambda$5875, where the absorption is stronger than the emission component. The presence of narrow He lines, identifiable only thanks to our high resolution, would imply a Type Ibn classification for SN 2024abvb, although we note that the weakness of these features could also lead to labelling it as a transitional Ibn/Icn SN (analogous to the transitional Ibn/IIn SNe; Pastorello et al., 2015). The possibility of hidden He in Icn SNe was briefly discussed in Gal-Yam et al. (2022) (Extended Data Figure 10). This suggests that an empirical classification based on low-resolution spectra may be unreliable, especially when interaction with a CSM is present and could have broader implications for the physical interpretation and classification of Icn SNe as a whole.
@@ 3.1.2 XShooter post peak spectra
C $i$ lines are also prominently detected, with all expected strong transitions clearly visible. These lines, like He i, display P Cygni profiles with a velocity of roughly 120 km s-1 as measured from the absorption component. In contrast, there is no evidence for the presence of He $i$, C iv, N $i$ or N i lines in the spectra. The presence of C iii and O iii remains uncertain; if present, these features are extremely weak and appear only in absorption. Nevertheless, these tentative identifications, along with the non-detection of higher ionisation lines, place constraints on the ionising radiation field, limiting it to a maximum of approximately 35–40 eV, with the bulk of ionisation radiation likely peaking in the 10–25 eV range (Moore, 1945). O i lines are also detected and are mostly attributable to recombination processes. [O i] $\lambda 6300$ is detected in emission at $\phi=17.2$d with a velocity width of the order of $\sim 100\penalty 10000\$km s-1, suggesting a CSM origin.
@@
Among other lines, Si $i$ is identified, with multiplets 1, 2, and 3 being present in the spectra. Multiplet 4 is absent, and the strongest line of multiplet 5, if present, appears extremely weak. Mg $i$ is observed, particularly in multiplets 4, 6, and possibly 8, which seem to appear solely in emission. Conversely, multiplets 2 and 10 are undetected, which is consistent with their known atomic parameters. Lines at 5696.6 and 5722.7 Å could be tentatively identified as Al iii (multiplet 2). Al $i$ (multiplets 1 and 2) is still detected as in the UVES spectrum. This identification excludes the possibility that the feature at 5696.0 Å is due to C iii (singlet transition). However, the line at 4325.7 Å remains unidentified, with multiple possible line assignments requiring further investigation. Notably, as in the FORS2 and classification spectra, none of the C iii triplet transitions are detected. Fe $i$ lines are observed prominently in the XShooter spectra, particularly from multiplet 42, as well as multiplets 49, 48, and 37. These lines show P Cygni profiles, at times complex due to proximity with other lines, displaying an absorption minimum velocity of 100 km s-1. By the time of the second XShooter spectrum ($\phi=17.2$d; Fig. 21), Fe $i$ features have significantly strengthened, suggesting the possible formation of an “iron blanket” in the near-ultraviolet region ($\leq 3500$Å) similar to the usual pseudo-continuum formed by Fe lines in SNe IIn (e.g. Smith, 2017, and references therein). In addition to the emission lines at 5889 Å and 6003 Å observed in the UVES spectrum, another emission line at 8235 Å is identified in the XShooter spectra. For this feature, potential associations include C $i$, Mg $i$ and O i. However, since its profile differs from those of other C $i$, Mg $i$ and O i lines, the precise nature remains uncertain. This feature may represent a blend of contributions from all three species.
@@
The ionisation potentials for Mg $i$ and Si $i$, assuming the observed lines arise from recombination, are approximately 15 and 16 eV, respectively. Their detection thus broadens the inferred ionisation range while slightly lowering the average ionising energy with respect to the analysis of the non-metal elements. This reinforces the conclusion that the source is dominated by relatively soft UV radiation.